An airplane takes off and flies 245 km on a course of 45 degrees south of west and then changes direction and flies 175 km due south, where it lands. Find the displacement from its starting point to its landing point.
Found 2 solutions by Alan3354, rmos1967:
Answer by Alan3354(69443)

An airplane takes off and flies 245 km on a course of 45 degrees south of west and then changes direction and flies 175 km due south, where it lands. Find the displacement from its starting point to its landing point.
----------
You have 2 sides and the included angle.
Use the Cosine Law.

Answer by rmos1967(1)

Ax=245cos(45)=173.24 km
Ay=245sin(45)=173.24 km
Bx=0
By=175 km
Rx=173.24 West
Ry=173.24 + 175= 348.24 South
√(173.24)^2+(348.24)^2= 389 km
tan-¹(384.24/173.24)=63.5° South of West